1.  A solution is made by dissolving 125 g Na2SO4 in 65.0 g water.  Calculate the mole fraction of the solute and the solvent.  (MM Na2SO4 =  142 g/mole;  water = 18 g/mole)
1
Expert's answer
2021-04-25T03:49:47-0400

2NaOH+H2​SO4​→Na2​SO4​+2H2​O

= 65/18 = 3.61moles

125/142 = 0.88M

Total moles = 4.91

H2O = 3.61/4.91

= 0.74

Na2SO4 = 0.88/4.91

= 0.179